OpenStudy (anonymous):

A 50-gallon barrel of milk is 6% butterfat. How much skim milk should be mixed to make milk that is 3% butterfat?

OpenStudy (e.cociuba):

let x="added milk" __ 50(6%)+x(0%)=(50+x)(3%) __ multiplying by 100 to clear percents 300+0=150+3x __ 150=3x __ 50=x